Two cars, A & B are 400 miles apart. They start to approach each other, car A at 50 miles per hour and car B at 30 miles per hour. A canary is on car A when they start out. It flies until it reaches car B, then flies back to A, then back to B and continues until the cars meet. If the canary flies at a rate of 60 miles per hour, how far will it fly until the cars meet?
Answer by ankor@dixie-net.com(15639)

Two cars, A & B are 400 miles apart. They start to approach each other, car A at 50 miles per hour and car B at 30 miles per hour.
A canary is on car A when they start out.
It flies until it reaches car B, then flies back to A, then back to B and continues until the cars meet.
If the canary flies at a rate of 60 miles per hour, how far will it fly until the cars meet?
:
Let t = travel time when the two cars collide
50t + 30t = 400
80t = 400
t = 400/80
t = 5 hrs
:
Find how far the canary flies at 60 mph for 5 hrs
5 * 60 = 300 miles